5. Explain why the set S is not a basis for R3 .

u1= (-1, 3, 2), u2= ( 6,1,1) for R^3.

6. Given matrix A in row cchelon form (R),
find the (a) basis for row space of A and rank (A).
(b) basis for the solution space of A and nullity (A).

1 -2 0 0 3
R= 0 1 3 2 0
0 0 1 1 0
0 0 0 0 0
1
Expert's answer
2019-12-23T09:04:38-0500

5. A set "S = (\bold e_1, ..., \bold e_n)" is a basis in linear space "V" if for any "\bold v \in V" exists a unique combination of coefficients "v_i, i = 1..n", such that "\bold v = \sum_{i=1}^n v_i \bold e_i". Hence, it suffices to show that some vector in "\R^3" can not be expressed as linear combination of "\bold u_1, \bold u_2". Choose, for instance vector "\bold u_3 = \bold u_1 \times \bold u_2 = (1, 13, -19)" (Geometric intuition: since "\bold u_1", "\bold u_2" span a plane in "\R^3", vector "\bold u_3", which is cross product of two, will be perpendicular to the plane, hence it is impossible to express it as linear combinations of them). Now, solve "\lambda_1 \bold u_1 + \lambda_2 \bold u_2 = \bold u_3", which is "\left(\begin{array}{cc|c} 
 -1 & 6 & 1\\ 
 3 & 1 & 13\\
2 & 1 & -19 
\end{array}\right)". Reducing the augmented matrix to RREF, obtain "\left(\begin{array}{cc|c} 
 1 & 0 & 0\\ 
 0 & 1 & 0\\
 0 & 0 & 1
\end{array}\right)".

Obviously, this system has no solutions, because of the last row. Hence, S is not a basis in "\R^3".

6.

a) Rank of the matrix is the number of non-zero rows in RREF of matrix, hence "rk (A) = 3".

Basis of the row space consists of three non-zero row-vectors of the RREF of the matrix: "\bold v_1 = (1, -2, 0, 0, 3), \bold v_2 = (0,1,3,2,0), \bold v_3 = (0,0,1,1,0)".

b) Since first three column vectors of RREF of the matrix are pivots, first three columns of the original matrix constitute the basis of solution space of "A".

According to rank-nullity theorem, "nullity (A) = 5 - rk (A) = 5 - 3 = 2".